# Free Descriptions of Convex Sets

### v2

@@ 1.3 Notation and Basics
[…]
@@
Symmetric algebra $&\mathrm{Sym}^{k}(\mathbb{V})&&=\mathrm{span}\{v_{1}\cdots v_{k}:v_{i}\in\mathbb{V}\}\\ &&&=\{\textrm{polynomials of degree }=k\textrm{ on }\mathbb{V}\}\\ &&&=\{\textrm{symmetric tensors of order }k\textrm{ over }\mathbb{V}\}\\ &\mathrm{Sym}^{\leq k}(\mathbb{V})&&=\bigoplus_{i=0}^{k}\mathrm{Sym}^{i}(\mathbb{V}).\end{aligned}$
Alternating algebra $&\bigwedge\nolimits^{k}\mathbb{V}&&=\mathrm{span}\{v_{1}\wedge\cdots\wedge v_{k}:v_{i}\in\mathbb{V}\}\\ &&&=\{\textrm{skew-symmetric tensors of order }k\textrm{ over }\mathbb{V}\}.\end{aligned}$
@@
Spaces of invariants $&\mathbb{V}^{\mathsf{G}}=\{v\in\mathbb{V}:g\cdot v=v\textrm{ for all }g\in\mathsf{G}\},\\ &\mathcal{L}(\mathbb{V},\mathbb{U})^{\mathsf{G}}=\{A\in\mathcal{L}(\mathbb{V},\mathbb{U}):gA=Ag\textrm{ for all }g\in\mathsf{G}\}.\end{aligned}$
@@
[…]
@@
together with the following action of $\mathsf{G}$: If $g\in\mathsf{G}$ is (uniquely) written as $gg_{i}=g_{j}h$ for some $i,j\in[k]$ and $h\in\mathsf{H}$, then $g\cdot g_{i}v=g_{j}(h\cdot v)$ for any $v\in\mathbb{V}$. This construction is independent of the choice of coset representatives.
[…]
@@ 2.2 Presentation Degree
    Fix $d\in\mathbb{N}$ and a $\mathsf{G}_{d}$-representation $\mathbb{W}$ on which $\mathsf{H}_{d,d}$ acts trivially. For each $n\geq d$, we view $\mathbb{W}$ as a $\mathsf{G}_{d}\mathsf{H}_{n,d}$-representation on which $\mathsf{H}_{n,d}$ acts trivially. The associated $\mathscr{V}$-induction sequence is defiend as:
@@
As $\mathscr{U}$ is presented in degree $k$, there exists an algebraically free $\mathscr{V}$-module =$\mathscr{F}=\{\mathbb{F}_{n}\}$ and a surjective morphism $\mathscr{F}\to\mathscr{U}$ such that both its kernel $\mathscr{K}=\{\mathbb{K}_{n}\}$ and $\mathscr{F}$ itself are generated in degree $k$. Because each map $\mathbb{F}_{n}\to\mathbb{U}_{n}$ is a $\mathsf{G}_{n}$-equivariant surjection with kernel $\mathbb{K}_{n}$, its restriction to invariants $\mathbb{F}_{n}^{\mathsf{G}_{n}}\to\mathbb{U}_{n}^{\mathsf{G}_{n}}$ is surjective with kernel $\mathbb{K}_{n}^{\mathsf{G}_{n}}$.
@@ 2.3 Constructions of Consistent Sequences
Fix a family of group $\mathscr{G}=\{\mathsf{G}_{n}\}_{n\in\mathbb{N}}$ such that $\mathsf{G}_{n}\subseteq\mathsf{G}_{n+1}$. Suppose $\mathscr{V}=\{(\mathbb{V}_{n},\varphi_{n})\}$ and $\mathscr{U}=\{(\mathbb{U}_{n},\psi_{n})\}$ are consistent sequences of $\mathscr{G}$-representations. Then the following are also consistent sequences of $\mathscr{G}$-representations:
  1. missingfnum@enumiitem(Sums)
@@
  2. missingfnum@enumiitem(Tensors)
@@
    If $\mathbb{W}$ is a fixed vector space, viewed as a trivial $\mathsf{G}_{n}$-representation for all $n$, denote $\mathscr{V}\otimes\mathbb{W}=\{(\mathbb{V}_{n}\otimes\mathbb{W},\varphi_{n}\otimes\mathrm{id}_{\mathbb{W}})\}$.
  3. missingfnum@enumiitem(Polynomials)
@@
  4. missingfnum@enumiitem(Moments)
@@
  5. missingfnum@enumiitem(Images and Kernels)
    If $\{A_{n}\in\mathcal{L}(\mathbb{V}_{n},\mathbb{U}_{n})^{\mathsf{G}_{n}}\}$ is a morphism mapping $\mathscr{V}\to\mathscr{U}$, then the images $\mathrm{Im}\mathscr{A}=\{(A_{n}(\mathbb{V}_{n}),\psi_{n})\}$ and kernels $\mathrm{ker}\mathscr{A}=\{(\ker A_{n},\varphi_{n})\}$ form consistent sequences.
@@
  1. missingfnum@enumiitem(Sums)
@@
  2. missingfnum@enumiitem(Images and kernels)
    If $\mathscr{A}\colon\mathscr{W}\to\mathscr{U}$ and $\mathscr{A}^{*}$ are morphisms, then $\mathrm{im}\mathscr{A}$ and $\ker\mathscr{A}$ are generated in degree $d_{W}$ and presented in degree $k_{W}$.
@@
  1. missingfnum@enumiitem(Tensors)
@@
  2. missingfnum@enumiitem(Sym and $\bigwedge$)
@@
In this case, we also have $\mathrm{Sym}^{2}\mathbb{R}^{n}=\mathbb{S}^{n}$ and $\bigwedge^{2}\mathbb{R}^{n}=\mathrm{Skew}(n)$. The space $\mathrm{Sym}^{2}(\mathrm{Sym}^{\leq k}\mathbb{R}^{n})$ can be viewed as symmetric matrices whose rows and columns are indexed by monomials of degree up to $k$ in $n$ variables. These sequences arises in optimization problems involving tensors, sums-of-squares polynomials, and moment sequences [64, 65].
@@
Note that the last two conclusions in Theorem 2.11 fail without the restriction to $\mathsf{G}_{n}=\mathrm{B}_{n},\mathrm{D}_{n}$, or $\mathrm{S}_{n}$. For example, if we endow $\mathscr{V}=\{\mathbb{R}^{n}\}$ with the standard actions of the cyclic groups $\mathsf{G}_{n}=\mathrm{Cyc}_{n}$ and use embeddings by zero-padding, then $\mathscr{V}$ is generated in degree 1 but $\mathscr{V}^{\otimes 2}$ is not finitely-generated since $\dim\mathcal{L}(\mathbb{R}^{n})^{\mathrm{Cyc}_{n}}=n$ does not stabilize.
@@ 2.4 Permutation Modules
    The projections $\mathcal{P}_{\mathbb{U}_{n}}\colon\mathbb{U}_{n+1}^{\mathsf{G}_{n+1}}\to\mathbb{U}_{n}^{\mathsf{G}_{n}}$ are isomorphisms for all $n\geq d$ if and only if (a) holds and the number of orbits in $\mathcal{A}_{n}$, which equals $\dim\mathbb{U}_{n}^{\mathsf{G}_{n}}$, is constant for all $n\geq d$.
@@
Here we say $\mathscr{U}$ agrees with a free module starting from degree $d$ if there is a free $\mathscr{V}$-module =$\mathscr{F}=\{\mathbb{F}_{n}\}$ and a morphism of sequences $\mathscr{F}\to\mathscr{U}$ such that $\mathbb{F}_{n}\to\mathbb{U}_{n}$ is an isomorphism for all $n\geq d$.
@@
    If $\hat{\mathcal{A}_{n}}\subseteq\mathcal{A}_{n}$ is a set of $\mathsf{G}_{n}$-orbit representatives, then $\mathbb{U}_{n}^{\mathsf{G}_{n}}\cong\mathbb{R}^{\hat{\mathcal{A}_{n}}}$ has a basis consisting of $\mathbbm{1}_{\alpha}^{(n)}=\sum_{g\in\mathsf{G}_{n}/\mathsf{Stab}_{\mathsf{G}_{n}}(\alpha)}e_{g\alpha}$ for $\alpha\in\hat{\mathcal{A}_{n}}$, where $\mathsf{Stab}_{\mathsf{G}_{n}}(\alpha)=\{g\in\mathsf{G}_{n}:g\cdot\alpha=\alpha\}$. Moreover, we have $\mathcal{P}_{\mathbb{U}_{n}}\mathbbm{1}_{\alpha}^{(n+1)}=\mathbbm{1}_{\alpha}^{(n)}$ for each $\alpha\in\mathcal{A}_{n}$. Thus, $\mathcal{P}_{\mathbb{U}_{n}}$ is an isomorphism between invariants if and only if orbit representatives for $\mathcal{A}_{n}$ are also representatives for $\mathcal{A}_{n+1}$.
@@ 2.7 Limits of Consistent Sequences
3 Structural Results on Free Descriptions
With the relevant background from representation stability in hand, we now proceed to prove the structural results discussed in Section 1.1.2 pertaining to freely-described convex sets. We leverage these results to provide a range of examples of free descriptions arising in various applications in Section 4. We begin with the following instructive example, which shows that freely-described sequences of convex sets need not satisfy either of our compatibility conditions. It also shows that a fixed convex set can extend to different freely-described sequences, depending on its conic description.

### v3

@@ 1.3 Notation and Basics
Even-signed symmetric group $\mathsf{D}_{n}=\{g\in B_{n}:g\textrm{ has an even number of }-1\textrm{ entries}\}$
@@
[…]
Alternating algebra $\begin{aligned} &\bigwedge\nolimits^{k}\mathbb{V}&&=\mathrm{span}\{v_{1}\wedge\cdots\wedge v_{k}:v_{i}\in\mathbb{V}\}\\ &&&=\{\textrm{skew-symmetric tensors of order }k\textrm{ over }\mathbb{V}\}.\end{aligned}$
@@
Spaces of invariants $\begin{aligned} &\mathbb{V}^{\mathsf{G}}=\{v\in\mathbb{V}:g\cdot v=v\textrm{ for all }g\in\mathsf{G}\},\\ &\mathcal{L}(\mathbb{V},\mathbb{U})^{\mathsf{G}}=\{A\in\mathcal{L}(\mathbb{V},\mathbb{U}):gA=Ag\textrm{ for all }g\in\mathsf{G}\}.\end{aligned}$
@@
We use $[n]$ to denote $\{1,\ldots,n\}$. For $i\leq j$ we denote by $(i,j)$ the transposition permuting letters $i$ and $j$. For real numbers $a<b$, we use $[a,b)$ to denote $\{x\in\mathbb{R}:a\leq x<b\}$. For $i\in[n]$, we denote by $e_{i}\in\mathbb{R}^{n}$ the $i$th standard basis vector with a 1 in the $i$th entry and zero everywhere else, and we write $e_{i}^{(n)}$ when we wish to emphasize the dimension. If $x\in\mathbb{R}^{n}$, we denote by $\mathrm{diag}(x)\in\mathbb{S}^{n}$ the diagonal matrix with $x$ on the diagonal. If $X\in\mathbb{R}^{n\times n}$, we denote by $\mathrm{diag}(X)\in\mathbb{R}^{n}$ the vector of its diagonal elements. All vector spaces in this paper are finite-dimensional real vector spaces equipped with an inner product $\langle\cdot,\cdot\rangle$ unless stated otherwise. We emphasize that some of the inner products we use are nonstandard, so the transpose of a matrix and the adjoint of the linear operator it represents may differ. Given a subspace $\mathbb{W}\subseteq\mathbb{V}$, we denote by $\mathcal{P}_{\mathbb{W}}\colon\mathbb{V}\to\mathbb{W}$ the orthogonal projection onto $\mathbb{W}$. We denote by $\mathbb{R}^{n}_{+}$ the cone of entrywise nonnegative length-$n$ vectors, and by $\mathbb{S}^{n}_{+}$ the cone of PSD $n\times n$ matrices. If $\mathbb{V}$ is a vector space, we let $\mathrm{Sym}^{2}_{+}(\mathbb{V})\cong\mathbb{S}^{\dim\mathbb{V}}_{+}$ denote the cone of PSD linear maps in $\mathcal{L}(\mathbb{V})$.
@@
together with the following action of $\mathsf{G}$: For each $g\in\mathsf{G}$ and $i\in[k]$, there is a unique $j\in[k]$ and $h\in\mathsf{H}$ satisfying $gg_{i}=g_{j}h$, and we let $g$ act on $g_{i}v$ for $v\in\mathbb{V}$ by $g\cdot g_{i}v=g_{j}(h\cdot v)$. This construction is independent of the choice of coset representatives.
As vector spaces, we have an isomorphism $\mathrm{Ind}_{\mathsf{H}}^{\mathsf{G}}(\mathbb{V})\cong\mathbb{V}^{|\mathsf{G}/\mathsf{H}|}$. Hence, an $\mathsf{H}$-invariant inner product $\langle\cdot,\cdot\rangle$ on $\mathbb{V}$ induces a $\mathsf{G}$-invariant inner product on $\mathbb{V}^{|\mathsf{G}/\mathsf{H}|}$ by setting $\langle g_{i}v,g_{j}u\rangle=\delta_{i,j}\langle v,u\rangle$ for $v,u\in\mathbb{V}$ and $i,j\in[|\mathsf{G}/\mathsf{H}|]$. Here $\delta_{i,j}=1$ if $i=j$ and zero otherwise. We have an isomorphism $(\mathrm{Ind}_{\mathsf{H}}^{\mathsf{G}}\mathbb{V})^{\mathsf{G}}\cong\mathbb{V}^{\mathsf{H}}$ sending $v\in\mathbb{V}^{\mathsf{H}}$ to $\sum_{i}g_{i}v\in(\mathrm{Ind}_{\mathsf{H}}^{\mathsf{G}}\mathbb{V})^{\mathsf{H}}$ and $\widetilde{v}\in(\mathrm{Ind}_{\mathsf{H}}^{\mathsf{G}}\mathbb{V})^{\mathsf{H}}$ to $\mathcal{P}_{\mathbb{V}}\widetilde{v}\in\mathbb{V}^{\mathsf{H}}$.
@@ 2.2 Presentation Degree
    Fix $d\in\mathbb{N}$ and a $\mathsf{G}_{d}$-representation $\mathbb{W}$ on which $\mathsf{H}_{d,d}$ acts trivially. For each $n\geq d$, we view $\mathbb{W}$ as a $\mathsf{G}_{d}\mathsf{H}_{n,d}$-representation on which $\mathsf{H}_{n,d}$ acts trivially. The associated $\mathscr{V}$-induction sequence is defined as:
@@
As $\mathscr{U}$ is presented in degree $k$, there exists an algebraically free $\mathscr{V}$-module $\mathscr{F}=\{\mathbb{F}_{n}\}$ and a surjective morphism $\mathscr{F}\to\mathscr{U}$ such that both its kernel $\mathscr{K}=\{\mathbb{K}_{n}\}$ and $\mathscr{F}$ itself are generated in degree $k$. Because each map $\mathbb{F}_{n}\to\mathbb{U}_{n}$ is a $\mathsf{G}_{n}$-equivariant surjection with kernel $\mathbb{K}_{n}$, its restriction to invariants $\mathbb{F}_{n}^{\mathsf{G}_{n}}\to\mathbb{U}_{n}^{\mathsf{G}_{n}}$ is surjective with kernel $\mathbb{K}_{n}^{\mathsf{G}_{n}}$.
@@ 2.3 Constructions of Consistent Sequences
Fix a family of groups $\mathscr{G}=\{\mathsf{G}_{n}\}_{n\in\mathbb{N}}$ such that $\mathsf{G}_{n}\subseteq\mathsf{G}_{n+1}$. Suppose $\mathscr{V}=\{(\mathbb{V}_{n},\varphi_{n})\}$ and $\mathscr{U}=\{(\mathbb{U}_{n},\psi_{n})\}$ are consistent sequences of $\mathscr{G}$-representations. Then the following are also consistent sequences of $\mathscr{G}$-representations:
  1. fnum@enumiitem(Sums)
@@
  2. fnum@enumiitem(Tensors)
@@
    If $\mathbb{W}$ is a fixed vector space, viewed as a trivial $\mathsf{G}_{n}$-representation for all $n$, denote $\mathscr{V}\otimes\mathbb{W}=\{(\mathbb{V}_{n}\otimes\mathbb{W},\varphi_{n}\otimes\mathrm{id}_{\mathbb{W}})\}$. Since $\mathbb{V}_{n}\otimes\mathbb{W}\cong\mathbb{V}_{n}^{\oplus\dim\mathbb{W}}$ as $\mathsf{G}_{n}$-representations, we can identify $\mathscr{V}\otimes\mathbb{W}$ with $\mathscr{V}^{\oplus\dim\mathbb{W}}$.
  3. fnum@enumiitem(Polynomials)
@@
  4. fnum@enumiitem(Moments)
@@
  5. fnum@enumiitem(Images and Kernels)
    If $\mathscr{A}=\{A_{n}\in\mathcal{L}(\mathbb{V}_{n},\mathbb{U}_{n})^{\mathsf{G}_{n}}\}$ is a morphism mapping $\mathscr{V}\to\mathscr{U}$, then the images $\mathrm{Im}\mathscr{A}=\{(A_{n}(\mathbb{V}_{n}),\psi_{n})\}$ and kernels $\mathrm{ker}\mathscr{A}=\{(\ker A_{n},\varphi_{n})\}$ form consistent sequences.
@@
  1. fnum@enumiitem(Sums)
@@
  2. fnum@enumiitem(Images and kernels)
    If $\mathscr{A}\colon\mathscr{W}\to\mathscr{U}$ is a morphism consisting of linear maps $\mathscr{A}=\{A_{n}\in\mathcal{L}(\mathbb{W}_{n},\mathbb{U}_{n})^{\mathsf{G}_{n}}\}$, and if the sequence of their adjoints $\mathscr{A}^{*}=\{A_{n}^{*}\in\mathcal{L}(\mathbb{U}_{n},\mathbb{W}_{n})^{\mathsf{G}_{n}}\}$ is a morphism $\mathscr{U}\to\mathscr{W}$, then $\mathrm{im}\mathscr{A}$ and $\ker\mathscr{A}$ are generated in degree $d_{W}$ and presented in degree $k_{W}$.
@@
  1. fnum@enumiitem(Tensors)
@@
  2. fnum@enumiitem(Sym and $\bigwedge$)
@@
Note that the last two conclusions in Theorem 2.11 fail without the restriction to $\mathsf{G}_{n}=\mathrm{B}_{n},\mathrm{D}_{n}$, or $\mathrm{S}_{n}$. For example, consider the sequence of cyclic groups $\mathsf{G}_{n}=\mathrm{Cyc}_{2^{n}}$ embedded as usual in each other. Let $\mathscr{V}=\{\mathbb{V}_{n}=\mathbb{R}^{2^{n}}\}$ be the consistent sequence with embeddings $\varphi_{n}(x)=x\otimes[1,0]^{\top}$ and the standard action of $\mathsf{G}_{n}=\mathrm{Cyc}_{2^{n}}$ by cyclically shifting coordinates. Then $\mathscr{V}$ is generated in degree 1 but $\mathscr{V}^{\otimes 2}$ is not finitely-generated since $\dim\mathcal{L}(\mathbb{R}^{2^{n}})^{\mathrm{Cyc}_{2^{n}}}=2^{n}$ does not stabilize.
@@ 2.4 Permutation Modules
    The projections $\mathcal{P}_{\mathbb{U}_{n}}\colon\mathbb{U}_{n+1}^{\mathsf{G}_{n+1}}\to\mathbb{U}_{n}^{\mathsf{G}_{n}}$ are isomorphisms for all $n\geq d$ if and only if either of the equivalent conditions in (a) holds and the number of orbits in $\mathcal{A}_{n}$, which equals $\dim\mathbb{U}_{n}^{\mathsf{G}_{n}}$, is constant for all $n\geq d$.
@@
Here we say $\mathscr{U}$ agrees with a free module starting from degree $d$ if there is a free $\mathscr{V}$-module $\mathscr{F}=\{\mathbb{F}_{n}\}$ and a morphism of sequences $\mathscr{F}\to\mathscr{U}$ such that $\mathbb{F}_{n}\to\mathbb{U}_{n}$ is an isomorphism for all $n\geq d$.
@@
    Let $\mathcal{O}_{1}^{(n)},\ldots,\mathcal{O}_{D(n)}^{(n)}$ be the $\mathsf{G}_{n}$-orbits in $\mathcal{A}_{n}$ and set $\mathcal{O}_{i}^{(n+1)}=\mathsf{G}_{n+1}\mathcal{O}_{i}^{(n)}$ for $i\in[D(n)]$. Note that $\mathbb{U}_{n}^{\mathsf{G}_{n}}\cong\mathbb{R}^{D(n)}$ has a basis consisting of orbit indicators $\mathbbm{1}_{\mathcal{O}_{i}^{(n)}}=\sum_{\alpha\in\mathcal{O}_{i}^{(n)}}e_{\alpha}$. Moreover, $\mathcal{P}_{\mathbb{U}_{n}}\mathbbm{1}_{\mathcal{O}_{i}^{(n+1)}}=\mathbbm{1}_{\mathcal{O}_{i}^{(n+1)}\cap\mathcal{A}_{n}}=\sum_{\alpha\in\mathcal{O}_{i}^{(n+1)}\cap\mathcal{A}_{n}}e_{\alpha}$. If $\mathcal{A}_{n+1}=\bigcup_{g\in\mathsf{G}_{n+1}}g\mathcal{A}_{n}$ and the number of orbits in $\mathcal{A}_{n}$ and $\mathcal{A}_{n+1}$ are equal, then $\mathcal{O}_{i}^{(n+1)}$ are precisely the orbits in $\mathcal{A}_{n+1}$. This implies that $\mathcal{O}_{i}^{(n+1)}\cap\mathcal{A}_{n}=\mathcal{O}_{i}^{(n)}$ for $i\in[D(n)]$, and hence thta $\mathcal{P}_{\mathbb{U}_{n}}$ is an isomorphism of invariants. Indeed, the inclusion $\supseteq$ is trivial while if $\alpha\in\left(\mathcal{O}_{i}^{(n+1)}\cap\mathcal{A}_{n}\right)\setminus\mathcal{O}_{i}^{(n)}$ then $\alpha\in\mathcal{O}_{j}^{(n)}$ for some $j\neq i$ and $\mathsf{G}_{n+1}\mathcal{O}_{i}^{(n)}=\mathsf{G}_{n+1}\mathcal{O}_{j}^{(n)}$ since these two $\mathsf{G}_{n+1}$-orbits are not disjoint (they both contain $\alpha$), and hence must be equal. This contradicts the fact that $\mathcal{O}_{i}^{(n+1)}$ are distinct. Conversely, suppose $\mathcal{P}_{\mathbb{U}_{n}}$ is an isomorphism of invariants. If $\alpha\in\mathcal{A}_{n+1}\setminus\bigcup_{g\in\mathsf{G}_{n+1}}g\mathcal{A}_{n}$ then $\mathsf{G}_{n+1}\alpha\cap\mathcal{A}_{n}=\emptyset$ and $\mathcal{P}_{\mathbb{U}_{n}}\mathbbm{1}_{\mathsf{G}_{n+1}\alpha}=0$, a contradiction. Therefore, we have $\mathcal{A}_{n+1}=\bigcup_{g\in\mathsf{G}_{n+1}}g\mathcal{A}_{n}$. Furthermore, the number of orbits in $\mathcal{A}_{n}$ and $\mathcal{A}_{n+1}$ are equal, since these are the dimensions of $\mathbb{U}_{n}^{\mathsf{G}_{n}}$ and $\mathbb{U}_{n+1}^{\mathsf{G}_{n+1}}$, respectively.
@@ 2.5 Uniform Representation Stability
More precisely, there is a standard indexing for irreducibles of $\mathsf{S}_{n},\mathsf{B}_{n}$, and $\mathsf{D}_{n}$ in terms of partitions of $n$, and a particular way to pad these partitions to form $\lambda[n]$, see [50, §2].
@@ 2.7 Limits of Consistent Sequences
3 Structural Results on Dimension-Free Descriptions
With the relevant background from representation stability in hand, we now proceed to prove the structural results discussed in Section 1.1.2 pertaining to freely-described convex sets. We leverage these results to provide a range of examples of dimension-free descriptions arising in various applications in Section 4. We begin with the following instructive example, which shows that freely-described sequences of convex sets need not satisfy either of our compatibility conditions. It also shows that a fixed convex set can extend to different freely-described sequences, depending on its conic description.